OpenStudy (anonymous):

What is the value of expression? 5−8 • 5−4 8 and 4 exponents @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you have \[\Large 5^{-8} * 5^{-4}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the bases are both 5 when you multiply two exponential expressions like this, with the same base, you add the exponents

jimthompson5910 (jim_thompson5910):

\[\Large 5^{-8} * 5^{-4} = 5^{-8 + (-4)} = ??\]

OpenStudy (anonymous):

32?

jimthompson5910 (jim_thompson5910):

how are you getting 32?

jimthompson5910 (jim_thompson5910):

the next step you need to do is to add the exponents -8 and -4 to get ???

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

-12

jimthompson5910 (jim_thompson5910):

\[\Large 5^{-8} * 5^{-4} = 5^{-8+(-4)} = 5^{-12}\] what is the last step?

OpenStudy (anonymous):

umm oh man idek

jimthompson5910 (jim_thompson5910):

what does a negative exponent tell you?

jimthompson5910 (jim_thompson5910):

think back to \(\Large 7^{-3}\)

OpenStudy (anonymous):

repeated divison?

jimthompson5910 (jim_thompson5910):

it means that you flip the fraction 5/1 to get 1/5 to get a positive exponent \[\Large 5^{-12} = \frac{1}{5^{12}}\]

jimthompson5910 (jim_thompson5910):

So, \[\Large 5^{-8} * 5^{-4} = \frac{1}{5^{12}}\] I'm not going to evaluate 5^12 since that number is so very big. So I'll just leave it as 5^12

OpenStudy (anonymous):

oooh okay